# Enhancing LLM’s Cognition via Structurization

### v1

@@ A.3 Human Evaluation
Payment. The payment for annotators is higher than the minimum wage in our country. Due to the anonymity limitation, we currently cannot provide the exact number.
@@ B.4 Additional Experiments for Hallucination Evaluation
In Sec. 4.2 in the manuscript, we mainly study structurization with LLaMA2 and GPT-3.5-Turbo on the AttrScore benchmark. For a thorough comparison, this section provides the investigation of employing another series of models, i.e., Qwen-7B and Qwen-72B [2], on AttrScore and FactScore datasets to further demonstrate our method’s efficacy. Results are presented in Sec. B.4 and Sec. B.4, respectively.
@@ Appendix D Broader Impacts
Appendix E Resource Requirement
We use 8 NVIDIA A100-80G GPUs to train our StruXGPT model, and leverage 1-2 NVIDIA A100-80G GPUs for all the inference experiments.
Appendix F Prompt Examples for Structurization
F.1 Prompt Templates

### v2

@@ A.3 Human Evaluation
Payment. The payment for annotators is 2,000¥ in total, which is higher than the minimum wage in our country.
@@ B.4 Additional Experiments for Hallucination Evaluation
Qwen models as the LLM evaluators. In Sec. 4.2 in the manuscript, we mainly study structurization with LLaMA2 and GPT-3.5-Turbo on the AttrScore benchmark. For a thorough comparison, this section provides the investigation of employing another series of models, i.e., Qwen-7B and Qwen-72B [2], on AttrScore and FactScore datasets to further demonstrate our method’s efficacy. Results are presented in Sec. B.4 and Sec. B.4, respectively.
@@
Incorporation with GPT-3.5/4 models and CoT techniques. In Sec. 4.2, we the popular chain-of-thought (CoT) technique for prompt augmentation on the advanced GPT-3.5-Turbo model on the AttrScore [66] benchmark, where Sec. B.4 is a simply copy of the experimental results. Our method presents comparable performance against CoT, and achieves better enhancement after integrating with CoT, illustrating the compatibility and extensibility to more advanced strategies.
Consequently, Sec. B.4 further presents the incorporation of our StruXGPT with the more powerful GPT-4 model and the CoT technique, showing consistent benefits for powerful GPT-3.5 and GPT-4 models, either with or without the CoT strategy. The experiments and discussions further validate the effectiveness of our StruXGPT approach.
Table A5: Integration with GPT-3.5 and CoT.
Method Attr. Contra. Extra. Avg.
- 72.0 30.4 71.7 58.0
+Ours 77.1 31.8 77.4 62.1
+CoT 76.4 35.3 74.4 62.0
+CoT+Ours 78.9 42.9 74.5 65.4
Table A6: Integration with GPT-4 and CoT.
Method Attr. Contra. Extra. Avg.
- 86.2 43.3 88.3 72.6
+Ours 87.6 48.3 89.7 75.2
+CoT 88.8 48.9 89.7 75.8
+CoT+Ours 88.5 52.8 90.3 77.2
B.5 Comparison with Other Augmentation Methods
Besides the summarization-based and prompting-based augmentation methods (in Sec. B.5 and Sec. B.4 respectively), here we also compare our structurization approach with AdaptLLM [7], which developed several domain-specific LLMs (in BioMedicine, Finance, and Low) via the proposed reading-comprehension training technique. The experimental results are presented in Tab. A7.
Table A7: Comparison with AdaptLLM on various domain benchmarks.
Domain Subset Metric Baseline AdaptLLM Ours
Medicine PubMedQA Acc 59.6 63.3 63.0
Finance ConvFinQA EM 29.2 41.5 36.5
Law SCOTUS mic-F1/mac-F1 28.3/10.8 30.0/17.8 30.6/15.6
General BoolQ Acc 55.7 53.9 58.2
Accordingly, our method can also boost the Llama-7b baseline for 3%-7% without training, while AdaptLLM requires costly continual training of the baseline model on each domain corpus. In particular, for the general reading comprehension task, AdaptLLM-Fin does not introduce significant boosts, while AdaptLLM-Bio/Law even cause performance drops, which is mainly because AdaptLLM’s domain-adaptation tuning will harm the general capability more or less. In contrast, our method does not alter the baseline model, but only structurizes the input context to enhance LLM’s cognition ability on downstream tasks, showing stable and consistent improvements (e.g., a 2.5% increase on BoolQ). Although our final performance is slightly inferior to the domain-specialized AdaptLLM, our generalizability emphasizes the contribution of our work, as we bring consistent enhancement across downstream domains and cause no degradation on general tasks.
@@ Appendix C Limitation and Future Work
Performance on General Evaluation Benchmarks As our method focuses on context structurization to enhance LLM’s cognition ability, for the tasks where no context is provided, our method may not bring significant enhancement and even get worse. The commonly used MMLU benchmark [21] is a typical scenario, where LLMs are asked to answer questions without context references but requiring their parametric knowledge (learned during large-scale pre-training), and context structurization does not help. If we insist on structurizing the question alone and feeding it into LLM’s inputs, the model may be disturbed by the introduced information from StruXGPT and generate wrong answers. As shown in Tab. A10, our method causes a slight 0.1% decrease (measured by OpenCompass protocol [^7] https://github.com/open-compass/opencompass) when taking LLaMA2-7B-Chat [53] as the baseline model.
Table A10: Evaluation on general benchmarks.
Model MMLU BBH
LLaMA2-7B-Chat 45.93 30.47
+Ours 45.84 31.30
On the other hand, we have also tested another common benchmark, BBH [49], which is designed to evaluate LLMs’ reasoning capability when dealing with several logical sentences/statements. In this case, our method can adapt well to highlight the logical relations and boost LLM’s reasoning abilities by 0.8%.
In conclusion, we suggest users apply structurization to long-form or logically complex contexts, while taking the original question as inputs when there is no context provided.
@@ Appendix D Broader Impacts
Appendix E Prompt Examples for Structurization
E.1 Prompt Templates